# Measuring the splashback feature: Dependence on halo properties and history

### v1

@@ 3.3 Splashback feature as a function of halo formation history
[…]
@@ 3.3.1 Concentration
Figure 8: Top: Depth of splashback features as a function of halo NFW profile concentration and binned by redshift shown by the colours. Middle: Width of splashback features as a function of halo NFW profile concentration with reshift shown by the colour. Bottom: Depth/Width of splashback features as a function of halo NFW profile concentration with redshift shown in colours.
Figure 9: At $z=0$, the concentration binned by halo mass shown with colour. The dots with error bars represent the results obtained by fitting Equation 7. The hollow triangles show the results obtained by substituting the scale radius $r_{s}$ in Equation 7 with that from the NFW profile. Top: Depth of splashback features as a function of halo concentration. Middle: Width of splashback features as a function of halo concentration. Bottom: Depth/width of splashback features as a function of halo concentration. The splashback features are more significant with lower concentration and higher halo mass. The similarity between the points and the triangles reveals the insensitivity of the splashback features to the fitting methods.
@@
We bin haloes at each redshift by concentration. By fitting the NFW profile to individual haloes, we extract the parameters $\rho_{s}$ and $r_{s}$ given in Equation 5. We calculate the concentration of individual haloes and select haloes with concentrations between $c=1$ and $c=19$ at intervals of 222. We then calculate the splashback features of the stacked set of haloes in each bin.
In Figure 8, we show the depth (top), the width (middle) and the depth-to-mass ratio (bottom) as a function of concentration for various redshifts shown by the colours. The depth decreases steadily with concentration while the width gradually increases with the concentration. The depth of haloes with various concentration and redshift is around $\mathcal{W}=1.8\sim 3.0$. The width initially increases quickly from $\mathcal{W}=0.50$ to $\mathcal{W}=2.5$ until concentration $c\sim 19$ and then decreases slightly from $c=2.5$ to $c=2.0$. We fit the splashback depth and width with the functions (depth: $\chi^{2}_{\nu}=2.5$; width: $\chi^{2}_{\nu}=2.07$):
@@
When taking the splashback depth-to-width ratio, in the bottom panel of the Figure 8, it consistently decreases with the concentration. The depth-to-width ratio decreases slowly with overall variation $\Delta\mathcal{D}/\mathcal{W}=3.0$. By taking the ratio between Equations 25 and 26, we apply the fittings to the splashback depth-to-width as a function of concentration and redshift
@@
The fitting function reveals that haloes with the same concentration have larger depth and width at present, while the splashback depth-to-mass ratio is smaller. This indicates that the splashback feature is more significant at higher redshift, given the same concentration. This is because, under fixed concentration, the inner density profile remains fixed. As a result, any variation in the splashback feature reflects changes in the outer halo dynamics, such as the halo mass accretion rate. Haloes at earlier times accrete mass more rapidly (Van Den Bosch, 2002; Zhao et al., 2003), and we observe that haloes at higher redshift exhibit smaller splashback widths and thus more distinct splashback features, as shown in Figure 5. Therefore, when halo profiles are varied primarily by their outer dynamics — as is the case when binning by concentration — splashback features become less significant at lower redshift. As a halo increases in the concentration, the depth decreases while the width increases. This matches the previous fittings. As a halo increases in concentration, the splashback features get less distinctive. This is because high-concentration haloes are typically older and have experienced slower accretion in recent times. In Reed et al. (2005), it was shown that haloes which form earlier develop steeper density profiles, as reflected by higher concentrations. This correlation arises because the inner structure of a halo is set during its early accretion phase and remains largely unchanged thereafter. Meanwhile, haloes with lower accretion rate tend to have higher concentrations. As we have seen in Section 3.2, slowly accreted haloes at the present have less distinct splashback features. The particles in such haloes are more dynamically mixed and have more time to virialise, which smooths out the sharp density drop at the splashback radius. In contrast, low-concentration haloes have recently acquired rapid mass accretion. This leads to a sharper, more pronounced splashback feature due to less mixing and more radial orbits (Diemer & Kravtsov, 2014; Adhikari, Dalal & Chamberlain, 2014; More, Diemer & Kravtsov, 2015).
@@
Furthermore, in order to verify that the splashback feature dependence on concentration is consistent with other results, we select haloes at present, bin them by halo mass $M_{200\text{m}}$ and concentration. The mass bins are the same as defined in Section 2.2.1. For each mass bin, we bootstrap haloes which are binned by concentration and extract the parameters in the NFW profile in Equation 5 and parameters in Equation 7 as well as the splashback features from the fitted density profiles. Then, by substituting the parameter $r_{s}$ measured from Equation 5 into Equation 7, we can simulate additional halo density profiles. Based on the new density profile, we extract the predicted splashback features. Therefore, we compare the measured splashback features (dots) with the predicted features (hollow triangles) in Figure 9. The measured splashback features match well with the predictions from the NFW profile, and the similarity between the points and the triangles in the Figure reveals the insensitivity of the splashback features to the fitting methods. This demonstrates that the observed concentration dependence of the splashback features—predicted using the NFW profile—is consistent with measured splashback features based on the DK14 model. The validation of the NFW-based fitting indicates that splashback features are strongly influenced by the inner halo profile, particularly through the scale radius.
@@ 3.3.2 Peak height
Figure 11: Comparison between the splashback width fitting functions. Dashed line represents the fitting in Equation 3.1 obtained by fitting the mass data directly ($\chi_{\nu}^{2}=2.2$) while the dotted line represents the fitting in Equation 31 derived from the relationship between the width and peak height ($\chi^{2}_{\nu}=1.2$).
The peak height of a halo provides a description of how large a density fluctuation the halo represents and is often used to relate halo mass to the underlying matter density field. It has a one-to-one correspondence with the mass of a halo and redshift.
Since we see redshift dependence in the relation between the depth and width of our splashback features with other properties such as mass, we now investigate whether this dependence persists with quantities like peak height that incorporate redshift in its definition. Diemer & Kravtsov (2014) found that the logarithmic slope profiles of dark matter haloes varies significantly with peak height. At $1\leq\nu\leq 1.5$, the logarithmic slope minimum reaches -3, while at $\nu\geq 3.5$, the minimum reaches -4. For all peak heights, the slope minimum is more negative at lower redshift. To quantify these changes, in Figure 10, we show the depth and width of the splashback feature as functions of the peak height and redshift of the halo.
@@
There is an increase of the depth with the peak height. At $z=0$, the depth increases from $\mathcal{D}=2.0$ to $\mathcal{D}=2.8$ while at $z=2$, the depth increases from $\mathcal{D}=2.2$ to $\mathcal{D}=2.7$ at $\nu<5$. This is consistent with the findings in Diemer & Kravtsov (2014) since the depth is more significant with larger peak height at lower redshift. However, there is significant scatter in the relationship between the depth of the splashback feature and the peak height. so the splashback depth has no strong dependence on peak height. We apply a single power law to fit the splashback depth ($\chi_{\nu}^{2}=1.1$):
@@
The splashback depth fitting function indicates that its dependence on peak height and redshift is uncorrelated.
The width, on the other hand, has a very strong dependence on the peak height. Above a peak height of $\nu\sim 1$, the width steadily decreases with peak height, with a width of $\mathcal{W}\sim 2.0$ near a peak height of $\nu\sim 1$ and a width of $\mathcal{W}\sim 1.0$ near a peak height of $\nu\sim 4$. The splashback width fitting function indicates that it is approximately linearly dependent on the peak height ($\chi_{\nu}^{2}=1.06$).
@@
[…]
There are two additional factors that may contribute to this trend. One is that haloes with larger peak height are in general more massive. As we saw in Section 2.2.1. Therefore, the positive correlation between the splashback depth and peak height and the negative dependence of splashback width on peak height is expected. Secondly, haloes with larger peak height have lower concentrations; the transition of the profile from $r^{-3}$ to $r^{-1}$ happens at larger radii, which makes the splashback width smaller as well at higher peak height.
@@
Overall, when composing the splashback depth and width, the depth-to-mass ratio demonstrates consistent dependence on the halo peak height with fitting functions ($\chi_{\nu}^{2}=0.72$):
@@
Equations 29, 30 and 32 indicate that by binning the haloes by peak height, the splashback feature dependence on peak height and redshift can be separated. The splashback features get more and more significant with increasing halo peak height and time, which is consistent with the findings of the splashback features dependence on halo mass.
@@ 3.3.3 Formation time
Figure 13: Haloes at $z=0$ are binned by the formation time in the left panel, and binned by the formation time and the halo mass in the right panels. The formation time is defined as the time the halo reaches half of its $z=0$ mass. Top: Depth of splashback features as a function of halo formation time. The dashed line represents the linear fitting while the dotted line represents the fitting derived from halo mass. Bottom: Width of splashback features as a function of halo formation time. The dashed line represents the logarithmic fitting while the dotted line represents the fitting derived from halo mass.
@@
In order to provide a quantitative measure of the splashback features on halo formation time, we define the formation time of a halo as the time when the halo reaches half of its current mass (at $z=0$). We select haloes at $z=0$ with formation time less than $z=1.5$ and bin the haloes with formation time by $\Delta z_{\text{form}}=0.2$. Above z=1.5, most haloes have not yet reached half their $z=0$ mass. As seen in Figure 14. In the left panel of Figure 13, the splashback depth decreases and the splashback width increases with an increase in halo formation time. The splashback depth decreases approximately linearly from 2.8 to 2.0 while the splashback width increases from 1.0 to 2.2 with a gradually slowed rate. It is shown that for haloes formed at $z_{\text{form}}<0.75$, the splashback width decreases at a relative faster rate comparing to the haloes formed at $z_{\text{form}}>0.75$.
We implemented the fitting functions of the depth and width of the splashback features to quantify the changes:
@@
with $\chi_{\nu}^{2}=0.45$. Both the depth and the width vary with halo formation time logarithmically. Therefore, the haloes formed earlier in time have less distinct halo boundaries thus shallower and wider splashback features. Older haloes tend to have smaller masses as they grow more slowly than massive haloes (McBride, Fakhouri & Ma, 2009; Zhao et al., 2009). We plot the mass distribution with halo formation time in Figure 14 to also verify this argument. It is shown that halo with lower formation time tends to have higher median mass. As the splashback depth increases with halo mass, it should decrease with halo formation time. This matches the observation in Figure 3.
Furthermore, as the splashback feature in terms of formation time is consistent with the result from halo mass in Section 2.2.1, we implement the mass model from Wechsler et al. (2002) and modify it such as the growth rate of halo mass is proportional to the halo current mass. Therefore, the halo mass fitting equation in terms of redshift is
@@
where $M_{0}$ is halo present mass $M_{200\text{m}}$ and $\alpha$ is a fitting parameter. In this model, the formation time is calculated as
@@
The best-fit parameters show strong degeneracy, with large amplitude $A$ and steep negative slope $B$, implying the model is not robust and well constrained by the data. The fit is shown as dotted lines in Figure 13. It is shown that both fitting functions of the splashback depth and the splashback width have good matching with the simulation data. The two fittings coincide between $0.1\leq z_{\text{form}}\leq 1.0$, and the fitting with mass model tends to have smaller depth and larger width at $z_{\text{form}}\geq 1.0$. This means the mass model predicts more distinct splashback features at relatively large $z_{\text{form}}$. Two fittings verify the consistency between the splashback mass and formation time dependence, and they are most accurate between $0.1\leq z_{\text{form}}\leq 1.0$. Within this region, the fittings are relatively reliable. Beyond that, which equations matches better with either simulations and observations is yet to know.
Taken together, the analyses of both halo concentration and formation time consistently point to a common physical interpretation: haloes that form earlier, and thus possess higher concentrations, tend to exhibit less distinct splashback features. This is likely because early-forming haloes have already established dense inner profiles, leaving little contrast between the virialised and infalling regions to generate sharp splashback signatures. In contrast, haloes that experience more recent growth accumulate matter predominantly in their outer regions, which affects less the splashback width. Therefore, the halo assembly history influences the splashback features by shaping the inner density profile of the halo. This is distinct from the results presented in previous sections where haloes at higher redshifts have more distinct splashback features. In this case, haloes at $z=0$ but that reached half their mass at earlier times have more distinct splashback features. correlated with halo splashback features.

### v2

@@ 3.3 Splashback feature as a function of halo formation history
Haloes are influenced by their formation history, including when and where a halo was formed. Shin & Diemer (2023) noted that older haloes show a higher value for $R_{\rm sp}$ by $\sim 3-5\%$ when controlled for accretion rate. Therefore, we investigate the splashback features dependence on the halo peak height, concentration as well as halo formation time respectively.
@@ 3.3.1 Concentration
Figure 8: Top: Depth of splashback features as a function of halo NFW profile concentration and binned by redshift shown by the colours. Middle: Width of splashback features as a function of halo NFW profile concentration with redshift shown by the colour. Bottom: Depth/Width of splashback features as a function of halo NFW profile concentration with redshift shown in colours.
Figure 9: At $z=0$, haloes are binned by mass and concentration (colour-coded). The dots with error bars show results from fitting the DK14 profile (Equation 7), while the hollow triangles show results obtained by substituting the scale radius $r_{s}$ in the DK14 profile with that from the NFW fit (Equation 5). Top: Depth of splashback features as a function of halo concentration. Middle: Width of splashback features as a function of halo concentration. Bottom: Depth/width of splashback features as a function of halo concentration. The splashback features are more significant with lower concentration and higher halo mass. The splashback features are more pronounced for haloes with lower concentration and higher mass. The close agreement between points and triangles demonstrates that these trends are insensitive to the specific definition of concentration, confirming the robustness of our conclusions to the fitting method.
@@
We bin haloes at each redshift by concentration. By fitting the NFW profile to individual haloes, we extract the parameters $\rho_{s}$ and $r_{s}$ given in Equation 5. We calculate the concentration of individual haloes and select haloes with concentrations between $c=1$ and $c=19$ at intervals of $2$. We then calculate the splashback features of the stacked set of haloes in each bin.
In Figure 8, we show the depth (top), the width (middle) and the depth-to-width ratio (bottom) as a function of concentration for various redshifts shown by the colours. The depth decreases steadily with concentration while the width gradually increases with the concentration. The depth of haloes with various concentrations and redshifts is around $\mathcal{D}=1.8\sim 3.0$. The width initially increases quickly from $\mathcal{W}=0.50$ to $\mathcal{W}=2.5$ until concentration $c\sim 19$ and then decreases slightly from $c=2.5$ to $c=2.0$. We fit the splashback depth and width with the functions (depth: $\chi^{2}_{\nu}=2.5$; width: $\chi^{2}_{\nu}=2.07$)
@@
When taking the splashback depth-to-width ratio, in the bottom panel of Figure 8, it consistently decreases with the concentration. The depth-to-width ratio decreases slowly with overall variation $\Delta\mathcal{D}/\mathcal{W}=3.0$. By taking the ratio between Equations 25 and 26, we apply the fittings to the splashback depth-to-width as a function of concentration and redshift
@@
The fitting function reveals that haloes with the same concentration have larger depth and width at present, while the splashback depth-to-width ratio is smaller. This indicates that the splashback feature is more significant at higher redshift, given the same concentration. This is because, under fixed concentration, the inner density profile remains fixed. As a result, any variation in the splashback feature reflects changes in the outer halo dynamics, such as the halo mass accretion rate. Haloes at earlier times accrete mass more rapidly (Van Den Bosch, 2002; Zhao et al., 2003), and we observe that haloes at higher redshift exhibit smaller splashback widths and thus more distinct splashback features, as shown in Figure 5. Therefore, when halo profiles are varied primarily by their outer dynamics — as is the case when binning by concentration — splashback features become less significant at lower redshift. As a halo increases in concentration, the depth decreases while the width increases. This matches the previous fittings. As a halo increases in concentration, the splashback features get less distinctive. This is because high-concentration haloes are typically older and have experienced slower accretion in recent times. In Reed et al. (2005), it was shown that haloes which form earlier develop steeper density profiles, as reflected by higher concentrations. This correlation arises because the inner structure of a halo is set during its early accretion phase and remains largely unchanged thereafter. Meanwhile, haloes with lower accretion rates tend to have higher concentrations. As we have seen in Section 3.2, slowly accreted haloes at the present have less distinct splashback features. The particles in such haloes are more dynamically mixed and have more time to virialise, which smooths out the sharp density drop at the splashback radius. In contrast, low-concentration haloes have recently acquired rapid mass accretion. This leads to a sharper, more pronounced splashback feature due to less mixing and more radial orbits (Diemer & Kravtsov, 2014; Adhikari, Dalal & Chamberlain, 2014; More, Diemer & Kravtsov, 2015).
@@
Furthermore, in order to verify that the splashback feature dependence on concentration is independent of the density profile fitting model chosen, we select haloes at present, bin them by halo mass $M_{200\text{m}}$ and concentration $c$ defined in Section 2.2.4, and fit them with different fitting models. Within each bin, we fit the halos with both the NFW profile in Equation 5 and the DK14 profile in Equation 7, and then extract the parameter $r_{s}$ which relates to the concentration. The $r_{s}$ in the NFW profile corresponds to the scale radius where the logarithmic slope of the density profile equals 2-2, whereas the $r_{s}$ in the DK14 profile is a more general shape parameter that sets the transition between the inner Einasto-like component and the steepening outer component and is coupled to additional parameters such as $\beta$ and $\gamma$. By substituting the parameter $r_{s}$ measured from Equation 5 with the $r_{s}$ in Equation 7, we can simulate an additional DK14 profile. Therefore, by comparing the measured splashback features from the DK14 profile (dots) with the additional DK14 profile with $r_{s}$ from the NFW profile (hollow triangles) in Figure 9, we observe that the two types of splashback features match well with each other. The similarity between the points and the triangles in the figure reveals the insensitivity of the splashback features to the exact definition of concentration and the density profile fitting model used. This demonstrates that the observed concentration dependence of the splashback features—predicted using the NFW profile—is consistent with measured splashback features based on the DK14 model. The validation of the NFW-based fitting indicates that splashback features are strongly influenced by the inner halo profile, particularly through the scale radius.
Figure 11: Comparison between the splashback width fitting functions. The dashed line represents the fitting in Equation 3.1 obtained by fitting the mass data directly ($\chi_{\nu}^{2}=2.2$) while the dotted line represents the fitting in Equation 31 derived from the relationship between the width and peak height ($\chi^{2}_{\nu}=1.2$).
@@ 3.3.2 Peak height
The peak height of a halo provides a description of how large a density fluctuation the halo represents and is often used to relate halo mass to the underlying matter density field. It has a one-to-one correspondence with the mass of a halo and redshift. Halos sharing the same $\nu$ correspond to equally rare initial density peaks, so they can be compared across redshift as systems at similar evolutionary stages and in similar environments.
Since we see redshift dependence in the relation between the depth and width of our splashback features with other properties such as mass, we now investigate whether this dependence persists with quantities like peak height that incorporate redshift in their definition. Diemer & Kravtsov (2014) found that the logarithmic slope profiles of dark matter haloes vary significantly with peak height. At $1\leq\nu\leq 1.5$, the logarithmic slope minimum reaches -3, while at $\nu\geq 3.5$, the minimum reaches -4. For all peak heights, the slope minimum is more negative at lower redshift. To quantify these changes, in Figure 10, we show the depth and width of the splashback feature as functions of the peak height and redshift of the halo.
@@
There is an increase of the depth with the peak height. At $z=0$, the depth increases from $\mathcal{D}=2.0$ to $\mathcal{D}=2.8$ while at $z=2$, the depth increases from $\mathcal{D}=2.2$ to $\mathcal{D}=2.7$ at $\nu<5$. This is consistent with the findings in Diemer & Kravtsov (2014) since the depth is more significant with larger peak height at lower redshift. However, there is significant scatter in the relationship between the depth of the splashback feature and the peak height, so the splashback depth has no strong dependence on peak height. We apply a single power law to fit the splashback depth ($\chi_{\nu}^{2}=1.1$)
@@
The fitting form in Equation 29 treats the dependencies on peak height and redshift as separable, i.e., no cross-terms are included. Thus the fitted parameters describe independent power-law scalings with $\nu$ and $z$, and no correlation between the two is assumed.
The width, on the other hand, has a very strong dependence on the peak height. Above a peak height of $\nu\sim 1$, the width steadily decreases with peak height, with a width of $\mathcal{W}\sim 2.0$ near a peak height of $\nu\sim 1$ and a width of $\mathcal{W}\sim 1.0$ near a peak height of $\nu\sim 4$. The splashback width fitting function indicates that it is approximately linearly dependent on the peak height ($\chi_{\nu}^{2}=1.06$)
@@
There are a few explanations for this relation. The higher width of the splashback feature at low $\nu$ is possibly due to the scatter in environments at low peak height. Some of these haloes may be located near massive neighbours in dense regions while others are isolated. This causes a higher variance in the stacked profiles near $R_{\rm sp}$ and causes the width to increase. Diemer & Kravtsov (2014) reported a much larger scatter for low-$\nu$ ($\nu<1$) haloes near massive neighbours compared to isolated ones, with $\frac{R_{\rm sp}}{R_{\rm 200m}}$ spreading across approximately $0.3$.
There are two additional factors that may contribute to this trend. One is that haloes with larger peak heights are in general more massive. As we saw in Section 2.2.1. Therefore, the positive correlation between the splashback depth and peak height and the negative dependence of splashback width on peak height are expected. Secondly, haloes with larger peak heights have lower concentrations; the transition of the profile from $r^{-3}$ to $r^{-1}$ happens at larger radii, which makes the splashback width smaller as well at higher peak heights.
@@
Overall, when composing the splashback depth and width, the depth-to-width ratio demonstrates consistent dependence on the halo peak height with fitting functions ($\chi_{\nu}^{2}=0.72$)
@@
Equations 29, 30 and 32 indicate that by binning the haloes by peak height, the splashback feature dependence on peak height and redshift can be separated. The splashback features get more and more significant with increasing halo peak height and time, which is consistent with the findings of the splashback features’ dependence on halo mass.
@@ 3.3.3 Formation time
Figure 13: Haloes at $z=0$ are binned by the formation time in the left panel and binned by the formation time and the halo mass in the right panels. The formation time is defined as the time the halo reaches half of its $z=0$ mass. Top: Depth of splashback features as a function of halo formation time. The dashed line represents the linear fitting while the dotted line represents the fitting derived from halo mass. Bottom: Width of splashback features as a function of halo formation time. The dashed line represents the logarithmic fitting while the dotted line represents the fitting derived from halo mass.
@@
In order to provide a quantitative measure of the splashback features on halo formation time, we define the formation time of a halo as the time when the halo reaches half of its current mass (at $z=0$). We select haloes at $z=0$ with formation time less than $z=1.5$ and bin the haloes with formation time by $\Delta z_{\text{form}}=0.2$. Above z=1.5, most haloes have not yet reached half their $z=0$ mass. As seen in Figure 14. In the left panel of Figure 13, the splashback depth decreases and the splashback width increases with an increase in halo formation time. The splashback depth decreases approximately linearly from 2.8 to 2.0 while the splashback width increases from 1.0 to 2.2 with a gradually slowed rate. It is shown that for haloes formed at $z_{\text{form}}<0.75$, the splashback width decreases at a relatively faster rate comparing to the haloes formed at $z_{\text{form}}>0.75$.
We implemented the fitting functions of the depth and width of the splashback features to quantify the changes
@@
with $\chi_{\nu}^{2}=0.45$. Both the depth and the width vary with halo formation time logarithmically. Therefore, the haloes formed earlier in time have less distinct halo boundaries and thus shallower and wider splashback features. Older haloes tend to have smaller masses as they grow more slowly than massive haloes (McBride, Fakhouri & Ma, 2009; Zhao et al., 2009). We plot the mass distribution with halo formation time in Figure 14 to also verify this argument. It is shown that haloes with lower formation time tend to have higher median mass. As the splashback depth increases with halo mass, it should decrease with halo formation time. This matches the observation in Figure 3.
Furthermore, as the splashback feature in terms of formation time is consistent with the result from halo mass in Section 2.2.1, we implement the mass model from Wechsler et al. (2002) and modify it such that the growth rate of halo mass is proportional to the halo current mass. Therefore, the halo mass fitting equation in terms of redshift is
@@
where $M_{0}$ is halo’s present mass $M_{200\text{m}}$ and $\alpha$ is a fitting parameter. In this model, the formation time is calculated as
@@
The best-fit parameters show strong degeneracy, with large amplitude $A$ and steep negative slope $B$, implying the model is not robust and well constrained by the data. The fit is shown as dotted lines in Figure 13. It is shown that both fitting functions of the splashback depth and the splashback width have good matching with the simulation data. The two fittings coincide between $0.1\leq z_{\text{form}}\leq 1.0$, and the fitting with the mass model tends to have smaller depth and larger width at $z_{\text{form}}\geq 1.0$. This means the mass model predicts more distinct splashback features at relatively large $z_{\text{form}}$. Two fittings verify the consistency between the splashback mass and formation time dependence, and they are most accurate between $0.1\leq z_{\text{form}}\leq 1.0$. Within this region, the fittings are relatively reliable. Beyond that, which equations match better with either simulations or observations is yet to be known.
Taken together, the analyses of both halo concentration and formation time consistently point to a common physical interpretation: haloes that form earlier, and thus possess higher concentrations, tend to exhibit less distinct splashback features. This is likely because early-forming haloes have already established dense inner profiles, leaving little contrast between the virialised and infalling regions to generate sharp splashback signatures. In contrast, haloes that experience more recent growth accumulate matter predominantly in their outer regions, which affects less the splashback width. Therefore, the halo assembly history influences the splashback features by shaping the inner density profile of the halo. This is distinct from the results presented in previous sections where haloes at higher redshifts have more distinct splashback features. In this case, haloes at $z=0$ but that reached half their mass at earlier times have more distinct splashback features.